Fermi National Laboratory


Accelerator Update - July 18-July 21

Friday July 18
The day shift began with Tevatron (TeV) and Recycler experts conducting studies, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 9:42 AM and at 11:41 AM established store 2800. Its initial luminosity was 30.74E30. Pbar resumed stacking.

The TeV suffered a sector D1 quench at 8:07 PM due to a quadrupole power supply (C:D0Q2) tripping off. Also, it seemed that the cryo system at C4 acted oddly.

Operators prepared the Collider Detector Facility (CDF) and D-zero (D0) for access into their collision halls at 8:45 PM, but only D0 accessed.

Operations put the TeV through a dry squeeze at 11:24 PM.

Saturday July 19
The midnight shift began with Operations preparing for a shot setup, and with MiniBooNE taking beam,

Operations put the TeV into shot setup at 1 AM. At 3:33 AM, Operations established store 2801 with an initial luminosity of 25.69E30. Pbar resumed stacking.

MiniBooNE tripped off at 8:18 AM for the first of five times over three and a half hours.

Pbar stacking stopped at 10:50 AM due to a kicker problem (A:IKIK). It was an hour before the device began working again.

Two Booster RF stations (BRF13 & 14) tripped off at due to a vacuum high-pressure alarm at 2:09 PM. The stations reset okay.

At 3 PM, A:IKIK faulted again; stacking stopped. Operations called in an expert. The kicker started working on its own at 3:31 PM. Stacking resumed. The expert arrived and made adjustments. The power amplifier tube for this device is slowly failing.

Two Booster RF stations (BRF5 & 6) tripped off due to a vacuum high-pressure alarm at 4:06 PM. The stations reset okay.

A Switchyard expert began a study period at 5:57 PM.

A Main Injector (MI) expert began a study period at 6:13 PM.

Two MI RF stations (MIRF 8 & 13) tripped off at 8:46 PM due to an overcurrent alarm. They both reset okay. But then the same stations tripped off three more times over the next hour and a half. Some of these trips may have been caused by the studies.

Operations terminated store 2801 at 10:14 PM. Operators put the TeV through a dry squeeze at 10:30 PM.

Sunday July 20
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 2803 at 1:45 AM. Its initial luminosity was 30.54E30. Pbar resumed stacking.

Switchyard experts began commissioning the SY120 line at 3 PM. By the time they had finished, they got beam as far as the target in the Meson train. (They now want to go to the next step and send beam to the MTest experiment area.)

MI experts conducted studies from 3:25 PM to 6:34 PM.

At 3:54 PM, A:IKIK again stopped working and halted stacking. Operators called in an expert, but the station started operating on its own at 4:30 PM. The expert made some adjustments at 4:43 PM and stacking resumed at 4:54 PM.

Operations terminated store 2803 at 7:04 PM and started the next shot setup at 7:40 PM.

Booster suffered a vacuum burst at 10:08 PM. Everything reset okay.

At 11:45 PM, Operations established store 2805 with an initial luminosity of 8.56E30. A thunderstorm pounded the lab just as this store was being established. A separator sparked several times due to lightning strikes. This stopped the luminosity from being all it could be.

Monday July 21
The midnight shift began with Operations monitoring store 2805, with Pbar stacking, and with MiniBooNE taking beam.

Operations terminated store 2805 at 1:27 AM. It was too low a luminosity to be of any use.

TeV experts began a beam maintenance study period at 2 AM.

All the MI RF stations tripped off at 2:16 AM due to a vacuum burst. Everything reset okay.

At 2:49 AM, a lightning strike caused a TeV anode power supply to trip, and a Pbar RF station (ARF1) and a MI RF station (MIRF5) to trip off. Everything reset except for MIRF5. The Tevatron Electron Lens (TEL) also tripped off along with a cryo compressor. Apparently there was a power glitch along with the lightning strike. Eventually, both the TEL and MIRF5 reset. The cryo compressor would not restart. It will be fixed later in the morning.

Operations loaded a 36x0 study store at 7:37 AM.

Today's Plan
The plan for this morning is to allow the experiments to access their collision halls. A shot setup will begin around noon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/21/2003   email Fermilab